HUBBLE SPACE TELESCOPE DAILY REPORT #2270 PERIOD COVERED: 0000Z (UTC) 12/14/98 - 0000Z (UTC) 12/15/98 Daily Status Report as of 349/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.2 Completed STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed WF/PC-2 7386 (A Snapshot Survey of Nova Shells) The WF/PC-2 was used to make observational images of the super nova shell Sgr-1 in order to detect and measure its expanding gaseous shell. These observations will trace the distribution of hot and cool gas in the nova shells, and show their degree of inhomogeneity on both large and small scales. The observations were completed as scheduled, and no problems were noted. 1.4 Completed Four Sets of NIC/1/2/3 8083 (FLATs: Warming Up - continuation) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the flat fields during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The observations for the third and fourth iterations of this proposal were not performed due to the NICMOS suspend event detailed in 2.3 which occurred near the beginning of the third iteration. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.5 Completed Three Sets of WF/PC-2 7377 (Dwarf Elliptical Galaxy Snapshot Survey II) The WF/PC-2 was used to make observations of the dwarf elliptical galaxy VCC1729 as part of Snapshot Survey II. This survey will provide important information on the globular cluster systems of such galaxies and the properties of their cores and nuclei. The observations were completed as planned, and no anomalies were reported. 1.6 Completed Six Sets of STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations taken in parallel with the prime science observations. The observations were completed as planned, and no problems were reported. 1.7 Completed Six Sets of WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed Eleven Sets of NIC/1/2/3 7963 (Darks: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The observations were completed as planned, and no anomalies were noted. 1.9 Completed WF/PC-2 6967 (Blue Continuum Snapshots of 3CR Radio - Continued Galaxies) The WF/PC-2 was used to make observations of the radio galaxies 3C180.0. Radio galaxies are an important class of extragalactic objects: they represent one of the most energetic astrophysical phenomena; they may be used as probes of their environments; and they are unique probes of the early Universe. The observations were completed as planned, and no problems were reported. 1.10 Completed Four Sets of STIS/MA1 7950 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA1) was used to make a dark calibration image. This is to establish if the detectors are performing nominally, that no internal movement of the MCP has taken place, and establish dark level in orbit for a given operating temperature. These darks will be used in later image reduction and also trended over instrument life time as indicators of detector degradation or damage. The observations were completed as planned, and no problems were reported. 1.11 Completed WF/PC-2 6754 (Imaging of the HH 30 Circumstellar Disk and Jets) The WF/PC-2 was used to make observations of the disk and jets of the Herbig-Harrow Object HH-30. Recent Hubble Space Telescope images have revealed HH-30 to be a prototype young stellar object system possessing an optically thick circumstellar absorption disk extended perpendicular to highly collimated bipolar jets. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed WF/PC-2 7387 (T-Tauri Star Snapshot Survey) The WF/PC-2 was used to make observations of the star DD-Tau as part of a T-Tauri class star survey. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed STIS/CCD 8069 (Throughput of F25ND5 Filter) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of star HD25340 in order to measure the throughput of the F25ND5 filter. Previous on orbit measurements of the throughput of this filter indicate that is about a factor of ~10 less efficient than the pre-launch estimates. The observations were completed as scheduled, and no anomalies were reported. 1.14 Completed STIS/CCD/MA1/MA2 7673 (STIS MAMA Sensitivity and Focus Monitor) The Space Telescope Imaging Spectrograph (CCD/MA1/MA2) was used to make observations of standard star GRW+70D582 in order to perform a MAMA sensitivity and focus monitoring program. This is to monitor sensitivity of each MAMA grating mode to detect any change due to contamination or other causes. The observations were executed as scheduled, and no anomalies were noted. 1.15 Completed WF/PC-2 6478 (The Evolution of Elliptical Galaxies in Distant Clusters) The WF/PC-2 was used to make observations in the V, R and I bands of the elliptical galaxies which are located in the distant galactic cluster Gal-Clus-0952+44. This will allow us to investigate the evolution of the cluster's ellipticals with redshift. The observations were completed as planned, and no problems were reported. 1.16 Completed NIC/2/3 7962 (NICMOS Warm-up Focus Monitoring) The Near Infrared Camera (Cameras Two and Three) was to be used to make observations of the globular cluster and nebula of NGC-3603 as part of the NICMOS warm-up focus monitoring. This proposal is used to determine and monitor the optimal focus settings for NIC/2 and NIC/3 during the dewar warm-up period at the end of 1998. The observations were not performed due to the NICMOS suspend event detailed in 2.3. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 9 Successful : 9 Scheduled Reacquisitions : 4 Successful : 4 2.2 FHST Updates: Scheduled : 27 Successful : 27 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared three times. As described in HSTAR 6815, NICMOS entered the suspend state at 348/154606Z when a status buffer message was received indicating that filter wheel #2 retried more than once to perform an operation. Per ROP NS-9, a NICMOS memory dump was performed at 348/1746Z and, per ROP NS-11, the NICMOS error counter was cleared at 349/0109Z. Preliminary analysis of the dump shows nothing unusual. Preliminary telemetry analysis shows no out of limit conditions on mechanism current or temperature. Further analysis continues. Proposals detailed in 1.4 and 1.16 above were affected by the suspend event. A status meeting is planned for this morning, when the filter wheel code will also be examined. NICMOS recovery and return to science operations are currently scheduled to begin on day 351. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH